I am trying to solve this equation:
3/x - 2/(x+1)=4/x
I know I need to get the LCD and that would be x(x+1).
I multiplied x+1 to the Num and dem of 3/x and I multiplied x to the num and dem of 2/(x+1). From there I get lost!

Answer by scott8148(5891)

good so far, but you also need to do the LCD on the 4/x ... that gives (3(x+1)/LCD)-(2x/LCD)=(4(x+1)/LCD)

multiplying by LCD gives 3(x+1)-2x=4(x+1) ... or 3x+3-2x=4x+4 ... -3x=1 ... x=-1/3